OpenStudy (anonymous):

help please!!!! The table below shows the values of f(n) for different values of n: n 1 2 3 4 5 6 f(n) 1 2 2 4 8 32 Which recursive function best represents the values shown in the table? (4 points)

OpenStudy (anonymous):

lalipapa123 help me pleasae

OpenStudy (anonymous):

somebody help please?

OpenStudy (lalipapa123):

wait for a while .i am trying to get it

OpenStudy (lalipapa123):

graph representation

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

these are the answers f(1) = 1, f(2) = 2, f(n) = f(n -1) f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -1) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) f(n - 2); n > 2

OpenStudy (lalipapa123):

did it helped u?

OpenStudy (anonymous):

mmm not really

OpenStudy (lalipapa123):

then?

OpenStudy (anonymous):

can u give me the answer please?

OpenStudy (lalipapa123):

which answer?

OpenStudy (anonymous):

of the question?

OpenStudy (lalipapa123):

this question ?

OpenStudy (anonymous):

yes

OpenStudy (lalipapa123):

i will try .just wait for a while

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

hey

OpenStudy (lalipapa123):

Try taking the ratio of successive terms and difference of successive terms when to find a recursive relation for a function. Or, try taking the difference of n th term and n-2 nd term or their ratio. Check a few combinations. Then see the values of the sequence obtained, if that matches with f(n), f(n-1), f(n-2) or their multiples. Some sequence will match. n = 1 2 3 4 5 6 f(n) = 1 2 2 4 8 32 f(n-1) = - 1 2 2 4 8 32 f(n-2) = - - 1 2 2 4 8 32 2 f (n-1) = - 2 4 4 8 16 f(n)-f(n-1) = - 1 0 2 4 24 f(n)-2 f(n-1) = - 0 -2 0 0 16 f(n) - f(n-2) = - - 1 2 6 28 f(n) / f(n-1) = - 2 1 2 2 4 f(n) / f(n-2) = - - 2 2 4 8 f(n) / f(n-3) = - - - 4 4 16 It seems to be that only the data items highlighted are matching. Now infer the relationship. f(n) / f(n-1) = f(n-2 f(n) = f(n-1) * f(n-2) , for 3 <= n <= 6 OR, f(n+2) = f(n+1) * f(n) , for 1 <= n <= 4 If we want the relation to be defined for the first two data items also, then we may define f(-1) = 1/2 and f(0) = 2, in that case, f (n) = f (n - 1) f (n - 2), for 1 <= n <= 6

OpenStudy (lalipapa123):

if it help you give me medal